### v1

# A Computational Intelligence Approach to Personal Scoring, Risk Stratification and Immune Age Characterisation from Common Blood Markers

Abstract
We introduce a family of indices based on a distance metric and closed learning algorithm for detecting rapid, small (precise) and personalised changes in blood test results. The indexes have the advantage of being fully explainable, unlike those featured in other statistical or machine-learning approaches. This metric takes the form of a quantitative measure or score that aggregates data from a set of numeric blood test results, combining and condensing them into a single numerical value that allows rapid triage, sorting, and immune age assessment. The main purpose of the score is to learn from, monitor, and provide a quick assessment of divergence from population and personalised reference values. We demonstrate the efficacy of the score against real and synthetic data from medically relevant cases, extreme cases, and empirical blood cell count data from the CDC NHANES survey spanning 13 years, from 2003 to 2016. We find that the score we introduce is informative when it comes to distinguishing healthy individuals from those with diseases, both self-reported and as manifest in abnormal blood test results, thereby providing an entry-level score for patient triaging. We show that the score varies over time and is correlated with biological age, leading to definition of an immune age as the inverse function of this relationship when taking different baskets of analytes based only on CBC test results, affording clinical evidence of its potential relevance to personalised precision medicine and individualised predictive healthcare.
Keywords: personalised medicine, quantitative precision medicine, pro-active healthcare, early detection, blood risk score, risk stratification, individual immune system assessment, immune age
1 Introduction
Risk assessment tools are often needed to improve patient safety and patient outcomes. Risk assessment scores for patient sorting and triaging are often qualitative, and are needed for the selection of patient pathways and health monitoring choices across all levels of health care.
Risk scores are valuable predictive tools that, when implemented, can be utilised by clinicians with positive results, leading to better patient outcomes. Risk assessment measures for patient triage are often needed to optimise patient pathways and health monitoring across all levels of health care. However, most scores are defined for very specific conditions, such as heart disease, cardiovascular disease (CVD) or venous thromboembolism (VTE), to mention just a few. Others reflect factors such as risk of severe disease [23, 31], clinical acuity [10, 33], and long-term outcomes [30, 32]. For example, the Intermountain risk score [19] provides a Complete/Full Blood Count (CBC/FBC) risk score shown to be associated with bleeding and in-hospital mortality based on variations in red blood cell size, such variations, originally used as an anaemia predictor [26], being often reported in FBC/CBC panels. FBC/CBCs have been found to contain enough information to yield clinically useful scores [3, 18]. At the same time, it has also been shown that although a CBC is a low-cost laboratory test that is almost universally used, its risk predictive information content is often underused [29].
Regular blood testing has been associated with short term mortality and with the ability to significantly improve emergency department triage [22]. Most risk scores depend on multiple factors external to a single test result in and of itself, and have proven useful when available, even in the context of the immune system [34]. A number of these scores are based on such factors as physiological data, medical and family history, and lifestyle choices, and reference a range of things, from age and sex to drinking and smoking habits. While numerical scores and colour-based scores are relatively common, to the authors’ knowledge no single general numerical and colour-based blood test risk assessment score exists that is based on a combination of distances from population or personalised average values, in particular for CBC and immune-related data.
In medical practice, a simple risk score based solely on blood test results and agnostic in regards to specific condition while still being able to quantify the results of a blood test, has not been proposed, to the authors’ knowledge. Here we introduce a general numerical and colour-based risk stratification score to quantify any number of abnormal blood markers. The risk model introduced is not data expensive when run ab initio, and yet it is shown to be informative, and its implementation is easily interpretable, while allowing for adaptation to enhance sophistication.
A marker, analyte or test parameter is a potentially non-mutually exclusive property related to a blood test. Analytes can include protein-based substances, antibodies, biochemical entities and any other product or sub-product of cellular function related to the blood or to the immune system. For example, in a Full (or Complete) Blood Count, there are usually about 13 to 15 analytes, although the proposed score is not limited to any particular set or number of such analytes. In the context of a cell’s properties, relevant tests can index, for example, cell count, cell size, cell morphology, cell nuclei morphology and cell maturity, among a wide range of parameters/analytes.
2 Methods
We call the risk assessment score introduced here the ”immune score”, given its connection to the immune system by way of the counts of blood cell subpopulations, in particular as registered in the most popular blood test, the Full Blood Count (FBC) or Complete Blood Count (CBC).
The immune score constitutes a dimensional reduction technique based on a single real-value number and a colour scheme that takes a multidimensional blood test space with dimension size equal to the number of analytes, where each value of a marker or analyte corresponds to the coordinate of that value in that dimension. The score itself can be seen as the norm of a suitable transformation of a vector pinpointing the health status of a patient in that space, relative to the test and set of markers or analytes. So the numerical vector value integrates all the analyte’s space dimensions.
Let
$\bar{x}=(x_{1},\dots,x_{n})$
be a vector of raw analyte values obtained by a suitable (series of) lab tests, as usually presented to physicians and patients.
We call the associated vector space for blood cell analytes the multidimensional immune space, or the immune space for short.
CDC NHANES data were used to estimate deficiencies and toxicities of specific nutrients in the population and subgroups in the U.S., to extract population reference data, and to estimate the contribution of diet, supplements, and other factors to whole blood levels of nutrients. This data can be used for research purposes and is publicly available [28].
2.1 Numerical score
This first formulation of the score takes the form of a linear function because it does not take into account possible interactions between analytes. We also assume that each analyte contributes equally to the immune score (this is likely to be only partially true, as some of the analytes are more or less medically informative than others, with this degree of informativeness itself varying with different conditions. Moreover, not all analytes are independent, and some analytes may be statistically dependent on others—suitable refinements will be introduced later).
For ease of comparison between successive versions of the score, all values are normalised within the range $[0,10]$. This means that each analyte will contribute with a maximum weight of
$w_{a}=\sqrt{\frac{100}{N}}.$
where $N$ is the number of analytes. From now on we will consider the example of $N=13$ typical for a CBC or FBC. From a geometrical point of view, a vector of values in the multidimensional immune space will be mapped to a point in a NND-“quadrant” going from 0 to $w_{a}$ that will be called the normalised immune space. This mapping will allow us to observe the diachronic evolution of an individual’s score as a trajectory in the normalised immune space where a notion of distance will come in handy.
While the score proposed is agnostic, in the sense that it can incorporate, or not, any number of analytes, and requires no other input, it can be adapted to incorporate medically relevant information by assigning weights that serve as multipliers of an analyte’s effect according to its criticality for different profiles.
The adoption of a score within a maximum range of $[0,10]$ can be approached in different ways. One is to assume no theoretical maximum values for analytes and asymptotically approach the maximum score of 10, without ever reaching it in individual cases. We adopted an alternative approach, capping values to a pre-established “reasonable” maximum, beyond which all specific values mean the same. The pre-established maximum is two standard deviations beyond normal (healthy) ranges, which sets the statistical significance critical value alpha at 0.05 (i.e., 95% confidence interval).
In an initial test, reference tables from Haematology Reference Ranges from the NHS were used for normal (healthy) ranges for each analyte according to age, sex, and pregnancy status [7]. $G$ will denote the set of possible categories:
[…]
newborn child, two-month-old child, six-month-old child,
one-year old child, 2-6-year-old child, 6-12-year-old child}\}
Let $g$ denote any value in $G$.
For each analyte, $r_{u}(g,i)$ is the upper limit in the normal range of the analyte $i$ for an individual belonging to group $g$. Similarly, $r_{l}(g,i)$ is the lower limit.
An efficient computer implementation of the score can pre-calculate the following values:
  • The expected vector, containing the mean value of each analyte for a given group:
    $\bar{e}(g)=(e(g,1),\dots,e(g,N)).$
    Mean values will derive from the analysis of suitable data to be obtained from reliable databases or directly collected by us. For ease of explanation, in the following we provisionally select the arithmetical average of the lower and upper limits in the NHS table:
    $e(g,i)=\frac{r_{l}(g,i)+r_{l}(g,i)}{2}.$
  • The standard deviation for an analyte $i$ in group $g$ is denoted by
    $\sigma(g,i).$
    Again, these values will come from a suitable set of cases. In the meantime, we will use the distance from the mean to either limit within the normal (healthy) range:
    $\sigma(g,i)=\frac{(r_{u}(g,i)-r_{l}(g,i))}{2}.$
  • The maximum difference vector, containing the maximum possible distance from the mean value of each analyte:
    $\bar{m}(g)=(m(g,1),\dots m(g,N)),$
    where
    $m(g,i)=\frac{(r_{u}(g,i)-r_{l}(g,i))}{2}+2\sigma(g,i).$
    (twice standard deviations from the limits of the healthy ranges).
  • The weighted vectors of each group (which will be used to normalise the possible values to a maximum immune score of 10):
    $\bar{w}(g)=(w(g,1),\dots,w(g,N)),$
    where $w(g,i)=w_{a}/m(g,i)$.
Let $\bar{c}=(c_{1},\dots,c_{N})$ be a vector of analyte values for an individual of group $g$. Then the normalised vector is calculated using the following formula:
$\bar{n}(\bar{c})=(f(c_{1},e(g,1),m(g,1),w(g,1)),\dots,f(c_{N},e(g,N),m(g,N),w(g,N))),$
where
$f(x,y,z,w)={\sf if}\>|x-y|\geq z\>{\sf then}\>w_{a}\>{\sf else}\>|x-y|\times w.$
The norm of the normalised vector will be called the Normalised Immune Score, or NIS for short:
$|\bar{n}(\bar{c})$$|$.
Further adding to the novelty impact of the metric, the score also offers a way to incorporate medical knowledge in the form of weights (scalar or piecewise function) to be incorporated natively, though non-linear analytes can also be introduced as synthetic analytes that represent or simulate such linearities, and would do so in a clear mathematical fashion rather than in an obfuscatory statistical fashion, making it more transparent than dealing with non-linearity directly or using other possible statistical black-box approaches. Unlike statistical or machine learning approaches, the distance-metric-based score is transparent, so that clinicians can fully understand how it works, and even reproduces the calculation by hand with a justification at every step without loss of generality.
2.2 Colour-based score
In addition to this numerical value, the score incorporates a second indicator that will be easy to read and interpret. The idea is to utilise colour codes as a means of flagging deviations from normal (healthy) reference values. This is called the Colour-Coded Immune Score or CCIS for short.
The CCIS uses the same analytes and their normal (healthy) ranges as before. There are thresholds for each of the analytes that trigger different types of alerts:
  • A value within normal healthy ranges: No alert, colour coded green.
  • An abnormal value within one standard deviation above/below healthy reference values: amber alert.
  • An abnormal value beyond one standard deviation from the normal healthy values: red alert. As in the normalised immune score, values are capped to a maximum of two standard deviations.
The results for each analyte are also combined in a single global value, which will constitute the CCIS. The rules are as follows:
  • Green: all values within normal range (green).
  • Amber: 1-3 individual amber alerts.
  • Red: More than 3 amber alerts or one or more red alerts.
Clearly, the CCIS is not a numerical value, but a colour-coded output, like the individual alerts for separate analytes. This clearly distinguishes the CCIS from the normalised immune score.
The information for the CCIS can be presented in a simple graphical way. We construct a doughnut graph with one slice for each analyte. Analytes assigned the same colour will be grouped together, which will produce doughnut graphs of at most 3 different colour-coded sections. Some examples for purposes of illustration are presented in Fig. 1.
3 A numerical score derived from the CCIS
We explored an alternative score that would follow the colour-coded scheme, but would also produce a numerical value. The intention was to combine the accessibility of the CCIS and the finer precision of the NIS in a single value. As with the latter, the value would range from 0 to 10.
We expected the following property: If $v_{1},v_{2}$ are possible score values and $v_{1}<v_{2}$, it should be the case that $v_{2}$ indicates greater cause for concern than $v_{1}$.
On the other hand, values should be closely related to the colour codes. That is, values $v_{1}$, $v_{2}$ and $v_{3}$ correspond to CCIS scores of green, amber and red, respectively, and they should be ordered as follows:
$v_{1}<v_{2}<v_{3}.$
A more precise way of expressing the above ideas is the set of rules codified in Table 3.
In the following section, we present a procedure for translating the CCIS into a numerical value according to the above schema. We will call this value the NCCIS. The expected behaviour of the NCCIS with respect to the number of analytes and its deviation from the norm can be observed in Fig. 3.
Table 1: NHS reference values: Typical tabular presentation of a Complete/Full Blood Count test.
Parameter/Component Result Low Ref Value High Ref Value
Haemoglobin 152.3 115 165
Total White Cell Count 9.3 3.6 11.00
Platelet count 286.5 140 400
Neutrophils 6.9 1.8 7.5
Lymphocytes 3.6 1.0 4.0
Monocytes 0.3 0.2 0.8
Eosinophils 0.2 0.1 0.4
Basophils 0.2 0.02 0.1
Red cell count 5.5 3.8 5.8
Haematocrit 0.4 0.37 0.47
Mean Cell Volume (MCV) 93.3 80 100
Table 2: Mean values and standard deviations found in the NHANES 2003–2016 database for the 13 analytes of the immune scores.
Analyte Mean Male Std Male Mean Female Std Female
Hb 150.72 11.51 133.33 11.62
WBC 6.92 3.12 7.01 2.09
Plt 228 54.33 255.42 64.19
RBC 4.93 0.45 4.40 0.38
MCV 89.96 5.13 89.22 5.72
Hct 0.44 0.03 0.39 0.03
MCH 26.96 10.33 26.55 10.47
MCHC 336.55 18.41 335.02 20.46
N 4 1.85 4.16 1.66
L 2.11 2.21 2.11 0.79
M 0.57 0.21 0.52 0.18
E 0.20 0.16 0.17 0.14
B 0.04 0.06 0.04 0.05
The behaviour of Fig. 3 illustrates how values at a remove from healthy values are pushed towards higher values by design, as an alerting mechanism.
3.1 Calculating CCIS and NCCIS
Some of the values in the calculation of the NIS:
  • $w_{a}=\sqrt{100/13}$ is the maximum normalised distance for each analyte in the normalised vector of a patient’s readings.
  • $e(g,i)$ is the mean (or expected) value for analyte $i$ in group $g$.
  • $\sigma(g,i)$ is the standard deviation from the mean.
  • $n(g,i)=(r_{u}(g,i)-r_{l}(g,i))/2$ is the normal (healthy reference) interval length from the mean value (assuming upper and lower limits are equidistant from the mean, an assumption to be revised in the future, both in the NIS and here in the NCCIS).
Now the raw (not normalised) borders between colours are
  • raw maximum value for green: $n(g,i)$.
  • raw maximum value for amber: $n(g,i)+\sigma(g,i)$.
  • raw maximum value for red: $n(g,i)+2\sigma(g,i)$.
And the normalised version:
$m_{gr}(g,i)$ $=$ $n(g,i)\times(w_{a}/(n(g,i)+2\sigma(g,i)))$
$m_{a}(g,i)$ $=$ $(n(g,i)+\sigma(g,i))\times(w_{a}/(n(g,i)+2\sigma(g,i)))$
$m_{r}(g,i)$ $=$ $(n(g,i)+2\sigma(g,i))\times(w_{a}/(n(g,i)+2\sigma(g,i)))=w_{a}$
Therefore the current NIS intervals for the different colours for individual analytes are:
  • Green: $0,m_{gr}(g,i)]$ (that is, from 0 distance from the mean value up to the normalised maximum value for green.)
  • Amber: $(m_{gr}(g,i),m_{a}(g,i)]$.
  • Red: $(m_{r}(g,i),w_{a}]$.
Figure 1: CCIS colour-based alerts. On the left, an individual with all values normal (healthy) and thus coded green. At the centre, an individual with two amber alerts. On the right-hand side, an individual with three red alerts, which produce a global red flag quantifying potential risk by colouring the numerical score depicted in the centre. In applications, colours following the guidelines of national health systems such as the NHS are recommended. For convenience, the numerical score can also be inverted to conform to the popular 0-10 grading scheme, where 10 is best and 0 is worst.
For global values the calculation has to take into account both the definition of the CCIS and the fact that each analyte can have proportionally different normal ranges and standard deviations:
  • Green: The minimum global green corresponds to 0 in every analyte. The maximum is $m_{gr}(g,i)$ in each analyte. This gives us the following interval:
    $0,\sqrt{\sum^{13}_{i=1}m_{gr}(g,i)^{2}}]$
  • Amber: The minimum global amber is 1 analyte value, just above green. The maximum is 3 maximum amber values and the rest (10) at the upper limit of green. Let analyte $j$ be such that $m_{gr}(g,j)$ is the minimum of the green upper limits, and let analytes $k$, $m$, $n$ be the three biggest of the amber upper limits. Then we have the following interval:
    $(m_{gr}(g,j),$
    $\quad\sqrt{(m_{a}(g,k)^{2}+m_{a}(g,m)^{2}+m_{a}(g,n)^{2})+m_{gr}(g,i_{1})^{2}+\cdots m_{gr}(g,i_{10})^{2}}].$
  • Red. The minimum global red is 1 red analyte value or 3 amber analyte values. The maximum is, obviously, 10 (13 maximum individual analyte values). Let us suppose that the lowest minimum for red is $(m_{r}(g,k))$ and that there are no 3 upper limits for amber analytes whose sum is below this. Then the intervals are:
    $(m_{r}(g,k),10].$
We will designate the ends of these intervals $\mbox{min}_{green}$, $\mbox{max}_{green}$, $\mbox{min}_{amber}$, $\mbox{max}_{amber}$, $\mbox{min}_{red}$ and $\mbox{max}_{red}$.
The NCCIS is calculated using the following function:
$\mbox{NCCIS}(v)$ $=$ ${\sf if}\>\mbox{CCIS}(v)=\mbox{green }\>{\sf then}\>a_{green}(\mbox{NIS}(v))$
$\>{\sf else}\>{\sf if}\>\mbox{CCIS}(v)=\mbox{amber }\>{\sf then}\>a_{amber}(\mbox{NIS}(v))$
$\qquad\>{\sf else}\>a_{red}(\mbox{NIS}(v))$
where CCIS and NIS are functions calculating the respective scores and the functions $g_{a}$, $a_{a}$ and $r_{a}$ map the NIS to the intervals set in the table at the beginning of Section 4:
$a_{green}(x)$ $=$ $w_{green}\times x$
$a_{amber}(x)$ $=$ $(10/3)+w_{amber}\times(x-\mbox{min}_{amber})$
$a_{red}(x)$ $=$ $(20/3)+w_{red}\times(x-\mbox{min}_{red})$
and the NIS value is weighted according to a normalising weight:
$w_{green}$ $=$ $10/(3\times\mbox{max}_{green})$
$w_{amber}$ $=$ $10/(3\times(\mbox{max}_{amber}-\mbox{min}_{amber}))$
$w_{red}$ $=$ $10/(3\times(10-\mbox{min}_{red}))$
Table 3: Mapping of score segments and colours. For convenience, the numerical score can also be inverted to conform to the popular 0-10 grading scheme, where 10 is best and 0 is worst.
Colour from CCIS Ideal interval for a numerical score
Green $[1,10/3]$
Amber $(10/3,20/3]$
Red $(20/3,10]$
3.2 NHANES Age Prediction
The zero counts were removed from the NHANES data prior to all analysis. However, the sex was replaced by 1 for males and 2 to denote females, for the quantitative analysis. The sex groups were separated by health status as was determined by binning the immune scores and ages by the healthy range thresholds as indicated on Table 1, which are as follows: RBC (between 3.8 to 5.8), and for the multivariate measure, the healthy ranges were as follows: normal MCH (27-33), HCT (37-47), lymphocytes (1.0-4.0), and neutrophils (1.8-7.5). Above, or below these healthy count ranges, the individuals were binned into the unhealthy groups. Further, the error rates analysis for the age prediction between the actual ages and immune-score computed ages were performed using these binned subsets.
4 Results
We have introduced a set of scores for patient triaging purposes tested against a large dataset based on numerical distance and a colour scheme. The most relevant part of the score is its single numerical value ranging from 0 to 10, capturing the distance of all other test values from population or personal reference values. Results show that such triage is statistically significant even in the face of noise, as in the NHANES database (containing 100K unfiltered surveys and tests), where disease is self-reported by interviewees who may have had the disease at the time of the survey or at some prior point in time. Filtering the score by test results also makes for a significant improvement, showing that it will perform better under less noisy conditions.
Figure 2: Immune score space: The $x$ axis represents the deviation of all the analytes from the expected value, while the $y$ axis measures the total number of analytes. A negative standard deviation indicates that the value is lower than expected and a positive one indicates a larger than expected value. The cut-off seen on the negative $x$ axis shows the fact that real life values for analytes have a strict lower bound but, technically, have no upper bounds. For instance, it is not possible to have a negative white blood cell count.
The behaviour of the numerical value is consistent through varying numbers of analytes and scales linearly with respect to deviation from the norm, and this graph explores the behaviour of the immune score over synthetically generated patient values as a function of number of analytes and divergence from healthy reference values.
Figure 3: The behaviour of NCCIS over synthetically generated patients with differing values and a deviation from the norm with respect to the respective expected values. The $x$ axis measures the stated analyte, while the $y$ axis measures the total number of analytes. In contrast to the behaviour of the immune score (Fig. 2), NCCISS is a non-linear function that displays a step-like behaviour by design, while remaining independent of the number of analytes.
Figure 4: Distribution of the Normalised Immune Score (NIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000–2018 database.The Healthy cohort is defined as those individuals with all analytes within the normal ranges as defined by the NHS (1). Others comprise the Unhealthy set. After filtering for adult status (18 years old or older) and completeness of analyte information, a total of 53,114 individuals are represented in the plot.
Figure 5: Distribution of the Normalised Colour-Coded Immune Score (NCCIS) from 0 to 10 ($y$ axis) among adult individuals in the NHANES 2000–2018 database. Individuals are defined as Healthy if all their analytes are within the normal ranges as defined by the NHS (1) and Unhealthy otherwise. After filtering for adult status (18 years old or older) and completeness of analyte information, a total of 53,114 individuals are represented in the plot.
In Table 4, we provide examples of CCIS and its correspondent NCCIS values for a sample of real-life and artificial cases. The second column of the table shows the NIS, for purposes of comparison. It is worth noting that these examples were generated using the NHS values for normal healthy subjects, and based on the provisional assumption that the distance between the upper and normal ranges is twice the standard deviation. The examples were taken from [8] as they were intended as very preliminary tests. We are aware that their source clearly states they are meant only for teaching purposes. Here they are used for purposes of illustration only, and the development of the NIS and NCCIS is not dependent on them.
One can observe how the NCCIS values meet the requirements we set in advance of their definition. We tested the NIS and NCCIS against real data from the National Health and Nutrition Examination Survey 2003–2016 (NHANES), provided in [28]. One measure of success was their ability to show how accurately the score discriminates between healthy subjects and patients suffering from various diagnosed illnesses.
To this end, we examined thousands of cases from [28] and calculated the distribution, mean values, and standard deviations for each of the analytes (see Fig. 7. We used the means and standard deviations from the data to define the metrics for both the NIS and NCCIS. Finally, we calculated the NIS and NCCIS for thousands of cases in the NHANES database to find individual scores. The resulting values were grouped as belonging to healthy individuals or to a selected list of common diseases.
In order to maximise the discriminatory power of both scores, we tried different alternative combinations of means, normal ranges and standard deviations, either taking these values from NHS values as of early 2022 or inferring them from the NHANES database (Fig. 7). At the end of the day, we settled for (1) NHS normal healthy reference values and means and standard deviations for the NIS; (2) means and standard deviations calculated from the NHANES database, and normal ranges from the NHS for the NCCIS. NHANES means and standard deviations are shown in Table 3.
Fig. 4 shows how NIS values are distributed according to different (self-declared) conditions for individuals in the NHANES database. It must be noted that the self-reported conditions were not independently confirmed. Moreover, the survey participants did not distinguish between current or past diagnoses. This could be expected to introduce some noise, as some currently healthy people will be labelled as having a condition and some people with conditions will not have been diagnosed. We hope to improve accuracy by filtering the data or by adding other data sources in the future.
Nevertheless, the NIS was able to discriminate between healthy and non-healthy individuals, as most healthy individuals are clustered around very low NIS values. In contrast, different conditions produced higher NIS values on average.
Conversely, the NCCIS did not substantially improve our knowledge of how some diseases impact analyte counts. There may be underlying co-morbidities or cases of blood-immune disorders that could create anomalous deviations from the predictions. Regardless, the methods are shown to be valid for predicting the complex immune profiles/trends (variation over time) observed in diverse patients when considering socio-clinical metadata parameters such as age, and sex, thereby suggesting their clinical applicability for purposes such as triaging and fast screening and extensions to such deviant clinical cases in prospective studies. In Fig. 5, it can be seen that reported medical conditions produce a higher average NCCIS value, but also induced some anomalous clustering. Furthermore, the distribution of scores was even wider than that of the NIS. We thus concluded that the NCCIS did enhance the knowledge already gleaned from the NIS and the CCIS.
Table 4: Different tested scores perform differently according to each score’s definition.
Individual NIS CCIS NCCIS
Adult male with all values at the mean 0.00 Green 0.00
Adult male values slightly removed from mean within normal healthy reference interval 1.84 Green 1.84
Adult male with all abnormal values 10.00 Red 10.00
Adult male with leukocytosis and diabetes 4.15 Red 7.60
Adult male with pancytopenia 5.54 Red 8.18
Adult male with mycosis 4.69 Red 7.83
Adult female with fatigue 3.07 Amber 5.43
Adult female with shortness of breath 6.70 Red 8.68
Adult female with thrombocytopenia 5.50 Red 8.16
Adult (female) with infection 3.63 Red 7.39
4.1 Linearisation and Separability
While the calculation of the score can incorporate normal ranges adjusted for age, race, sex, pregnancy stage, geography and any other consideration warranted by the literature, the score is intended to go beyond this clinical utility by capturing medical knowledge. A feature of the proposed (immune) score is the option to incorporate weights as multipliers in the form of scalars or piece-wise functions per analyte to modify its contribution relative to other analytes in a non-linear fashion, even under time sensitive conditions [11]. For example, in a blood differential test, conditions related to decreased white cell counts are milder than those associated with higher cell counts. However, the literature on conditions where there is a decrease of basophils, mast cells, monocytes and eosinophils in isolation is sparse, and reduced weights can be assigned to these possibly less-relevant markers.
Differences in cell shape and size, however, are usually more clinically significant and can be assigned greater weights. Typically, disorders affecting bone marrow function (i.e., blood cancers) result in the presence of abnormal (often immature) cells in peripheral blood; their presence in peripheral blood beyond this level would almost always be abnormal. A related point is whether the score can reflect subtle/minor changes in shape (variation in cell size, nuclear size, presence of other organelles) or only very crude and major deviations in size. For white blood cells, we do not yet understand what significance these have.
For example, the neutrophil-to-lymphocyte ratio relation is a key determinant of severity for sepsis, and involves 2 analytes. The synthetic analyte that can be added is the ratio itself, thus replacing a non-linear rule that would eventually make the score’s description too convoluted to read with ease with a key marker as another analyte [20, 25].
The scores introduced here are not intended to be used as a diagnostic tool on their own and can only quantify abnormality in terms of deviation from healthy reference (absolute or adaptable) values. The score is sensitive to out-of-range values and increases its value or changes its colour as a function of how removed values are from lower and upper bounds of reference values according to the number of standard deviations from the medians, but it cannot quantify diseases or conditions. The score indicates how far the bulk of all markers are from normal (healthy) reference values, the median, and an interval determined by published reference values for specific demographic or health conditions. Medians can be derived from empirical data (Fig. 6).
Figure 6: Distributions within ‘healthy’ reference values are heavily skewed in most cases, do not follow a normal or uniform distribution and often have long tails suggesting a personalised approach to reference values for abnormal results. The list of analyte labels is in the Supplementary Information.
Figure 12 shows how male and female NHANES data can be distinguished by the scores using Red Blood Cell (RBC) counts as a health status predictor. The normal versus abnormal RBC ranges from the NHS were used to label the person as healthy or unhealthy. The Pearson $R^{2}$ value for RBC-dependent health status was found to be 0.176, for both sexes considered together. The Pearson $R^{2}$ values were 0.203 and 0.455, for male and female RBC counts considered separately, when comparing healthy and unhealthy groups.
We observed that the self-reported health status of the patients did not accurately predict their clinically assessed blood analytes. This was expected because self-reports may have reflected illness at various points in time in a person’s life rather than ongoing conditions. In other words, with this experiment, we de-noised the data by disregarding self-reported status to test the score against cleansed data based exclusively on CBC results as a discriminant, according to medical guidance on normal population reference values. That is, the interpretation of the health status a person could have potentially received from a medical doctor based exclusively on their CBC.
When healthy and unhealthy patients were separated using a subset of four CBC/FBC analyte measures, namely, MCH, HCT, neutrophil, and lymphocyte counts, as shown in Figure 13, we obtained the second best $R^{2}$ value of 0.075, for both sexes. The $R^{2}$values were 0.050 and 0.123, for the male and female groups, respectively. We found that neutrophil count alone, as a single clinical parameter, had a near equal performance as the composite measure of these four analyte parameters, with an $R^{2}$ of 0.072 when including both sexes. The inclusion of other white blood cell (WBC) parameters, such as monocyte count, eosinophil, and basophil, did not outperform these correlates. We observed that different subsets of a CBC/FBC by health status separated the data. All reported measures were statistically significant, with a two-tailed P<0.0001<0.0001.
This also suggests that the mutlivariate biopsychosocial determinants of health should also be integrated for a top-bottom analysis of the disease predisposition and health dynamics of patients when assessing such self-reported measures. The irreducibility (multidimensionality) of these self-reported health measures may help explain why they are highly noisy and intersect with multiple other health factors, including multi-scale stressors such as the time of day, basic survival factors such as nutritional status/intake, sleep-related mood, psychological state, and social determinants of health, together with other health parameters.
4.2 Adaptive reference values
The reference ranges found in the medical literature are based on population-wide statistics that often fail to account for individual differences [6, 5]. To address this, we introduced an adaptive version of the immune score with the objective of providing personalised reference ranges to an individual’s medical history data. Our rationale is that, in applying the immune score to the resulting adaptive ranges, we obtain an immune score that is more significant in the statistical and clinical sense.
The main underlying assumption of this adaptive immune score definition is that, if a significant event (in the clinical sense) can be measured via an analyte, such an event will present itself as a value that breaks from the trend defined by the behaviour of the analyte over time. A second assumption is that such events can occur within the normal range as defined by a population reference range. A third assumption is that, while a trend becomes more incontrovertible the more data points there are, past values are less significant than proximate events. The next assumption is that the trend can be non-monotonous. The final assumption is that standard reference ranges can detect such events with acute sensitivity, while low specificity can thus be improved by narrowing the ranges but not by expanding them.
The adaptive feature of the immune score starts with a mathematical function that models the behaviour of an analyte as a time series (observed values over time). For each given point, the function assigns a linear approximation to the current and past data over a fixed time span (called a time step). The differential of this approximation is weighed against the differentials obtained for previous data via a linear combination of the current and previous derivatives. Finally, this linear approximation is evaluated over the mean time for the corresponding time-step. The resulting model is a function that carries the momentum of all previous values, and only previous values, yet remains adaptable for future long-term trends while being robust in the face of temporal disruptions to the trend. The speed of this adaptation can be controlled by hyper-parameters such as the size of the time step and the coefficients for the linear combination of differentials.
The function described in the preceding paragraph represents the adaptive mean, from which the adaptive ranges are inferred. The adaptive ranges consist of two functions, the upper limit and the lower limit. The upper limit is defined as the adaptive mean plus the standard deviation of all the values that are above the mean, up to the point of time in question. Similarly, the lower limit is the adaptive mean minus the standard deviation of all the values that are below the adaptive mean up to the relevant point in time. Next, we smooth the transition from the population ranges to the adaptive ranges by means of a linear combination of both ranges such that the adaptive range becomes the dominant factor in the time function.
Furthermore, this linear combination ensures that the adaptive ranges stay within the population ranges by defaulting to the population ranges whenever the resulting adaptive maximum is above the population maximum or the adaptive minimum is below the adaptive mean.
Figure 7: Statistical inference without smart outlier detection. In contrast to the smart outlier detection technology, simple statistical regression towards the observed values results in skewed inferred ranges.
Figure 8: Smart outlier detection example. The machine learning technology can adapt the range of expected values for all analytes, showing stable trends, while also detecting abnormal values or trends. In the image, we can show the result of machine learning inference of the ranges of normal values for a patient while also detecting abnormal values.
5 Immune age
The immune system plays an important role in protecting against infections and in the maintenance of human health. Blood is the medium in which most immune cells circulate through the body.
Ageing is associated with complex changes in and dysregulation of cellular processes. Nine tentative hallmarks that represent common denominators of ageing in different organisms have been described [24].
Worldwide, as lifespans are generally increasing, there is a concomitant increase in age-related morbidities, particularly cardiovascular diseases, and increased susceptibility to seasonal infections, cancers, and neurodegenerative disorders. A unifying link between the higher rates of these disparate diseases observed in ageing populations is the progressive decline in immune functions.
It has been postulated that inflammation plays a critical role in regulating physiological ageing. Inflammatory components of the immune system are often chronically elevated in aged individuals, a phenomenon that has been termed “inflammageing” [12, 14, 13, 15, 17, 2].
However, the dynamics of this process at the individual level have not been characterised, hindering quantification of an individual’s ‘immune age’.
[2] used multiple ‘omics’ technologies to capture population- and individual-level changes in the human immune system in individuals of different ages sampled longitudinally over a nine-year period. They observed high inter-individual variability in the rates of change of cellular frequencies that was dictated by the individual’s baseline values, allowing identification of steady-state levels toward which a cell subset converged and the ordered convergence of multiple cell subsets toward older adult homeostasis.
In [16], an IMM-AGE score was described that captured an individual’s immune-ageing process. The IMM-AGE score correlated with age, yet it captured additional metrics such as cell-cytokine response better than chronological age.
Here we present a method where a single blood test, as in a complete full blood count of 13 blood analytes, gives insight into the immune age as compared to biological age.
5.1 Immune Score Variation Over time
We hypothesised that as the human body ages the immune score will vary over time. In particular, we expected to see an upwards trajectory in direct relation to age.
To test the hypothesis, we selected two cohorts comprising all the entries with ages ranging between 20 to 30 years, separating them by sex (male or female). These cohorts were used to build two immune spaces by computing the respective mean and standard deviations for each of the thirteen analytes. Afterward, for each sex we selected seven cohorts of adults aged 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75, and 75 to 85 years of age, making for a total of 14 cohorts. Finally, we computed the immune score for each entry. The mean immune age for each of the cohorts is shown in Fig. 9.
Figure 9: (a) The mean immune age as a function of age for NHANES data. The reference ranges were computed over the population aged 20 to 30 for males and females, respectively. Each of the data points represents the mean immune score for the cohort [age5,age+5][age-5,age+5]; for example, 50 represents the immune score for the population aged 45 to 55 years. (b)The distribution of immune age as a function of age. While a correlation trend is visible especially for males, across sex weak because healthy and non-healthy cohorts (both self-reported and ground truth) are confounded.
The results show significantly different behaviour over time for male and female cohorts. For subjects over the age of 30, it is evident that the male population presents an upward trajectory with respect to age for the mean immune score, as well as an increase in the 50th percentile and variance. However, the trend in females is slowed down (compared to males) between ages 40 and 50 and reversed between ages 50 and 60. It remains to be investigated whether some of the analytes involved are confounded with contradictory trends from other physiological differences, such as menopause, which occurs exactly between the ages where the slow-down and reversal in the age score trend for females is detected. It is known and has been reported that pregnancy and menopause produce a higher and more progressive increase in red cell counts, haematocrit levels and increased mean cell volume (MCV), and haemoglobin concentrations [21, 4, 9, 27, 1].
5.2 Immune age based on immune score
We define the immune age as the reverse function of the relation between the expected immune score and the age of an individual. Formally:
Definition 5.1.
For a given immune score $\alpha$, the immune age $A:\alpha\mapsto a$ is defined as the function:
$A(\alpha)=a.\overline{\{S(p):p\text{ age is }a\})}=\alpha,$
where $\overline{\{S(p):p\text{ age is }a\})}$ is the mean immune score for all the individuals of age $a$.
In other words, the immune age is defined as the age in the function of the expected immune score; the immune age for a given score is the age cohort for which such a score is expected.
Given the observed behaviour of the mean immune score over time, we assert that we can apply the given immune age definition for males aged 20 or older. However, we have shown that the functional relation is much weaker, or nonexistent, in the female population.
As explained in Section 5.1, the data compiled comprises a list of seven values representing the mean immune score for each of the seven age cohorts [^1] The corresponding ages are 20 to 25, 25 to 35, 35 to 45, 45 to 55, 55 to 65, 65 to 75 and 75 to 85.. In order to apply the Definition 5.1, we fitted a linear model of degree 3 to the data, obtaining a monotonously increasing function that allows us to interpolate and extrapolate the expected immune score over specific ages. The resulting curve, which we call an immune age curve, is shown in Fig. 10.
Figure 10: A linear approximation of degree 3 to the mean score over seven age cohorts. The linear model smooths the curve and allows for extrapolation.
We will only consider integer ages in years. Therefore we can represent the immune age curve by a list of 70 real values that contains the inferred immune score for each age between 20 and 90. Let us denote this curve by the list of pairs $\text{IAC}=\langle(20,3.0),(21,3.1),\dots,(a,\alpha),\dots,(90,4.75)\rangle$.
We now can obtain the immune age $A$ by means of the following function:
$A(\alpha)=\begin{cases}20\quad&\text{if}\,\alpha\leq 3\\ 90\quad&\text{if}\,\alpha\geq 4.75\\ a.(a,\alpha^{*})\in\text{IAC}\quad&\text{elsewhere}\end{cases}$
where $\alpha^{*}$ is the closest score in $IAC$ to $\alpha$. If there are two scores at the same distance, we choose the leftmost option. The resulting function is shown in Fig. 13. The curve shows some minor bumps which are the results of the discretisation used. We maintain that this distortion is not significant.
Figure 11: The immune age is the age as a function of the expected immune score. The slight distortions are an artifact of the discretisation.
As a tool that returns an estimated age with respect to a score, we can analyse its predictive power. Using the NHANES data and focusing on males aged 20 to 90, we can measure the error of using the score to predict a person’s biological age.
As shown in Figure 14, prediction error rates are much lower in the healthy NHANES data, in comparison to the unhealthy groups, for both sexes, and can be therefore used to quantify healthy patient ages to some extent. The prediction error rates for the immune-score age, was overestimated roughly within 7 years, for the healthy groups, in both sexes. Thus, the error rate of age prediction for healthy people is small, with a mean error rate of 7.33±1.21 for females, and 7.62 ±1.17 for healthy males. Further, there is a statistically significant difference between the healthy and non-healthy groups (regardless of sex) in the error rates of age prediction. The Pearson correlation statistics on the predicted age error rates, by RBC count health status, as discussed before, was an R-square value of 0.293, and 0.168, for males (P<< 0.0001) and females (P<<0.013). Age prediction error rates were high and non-significant when the self-reported health measures were used to separate the health status groups but were significant when taking abnormal CBC values only.
Figure 12: Age Prediction Error Rates between actual patient (biological) age and Immune-score predicted age from NHANES dataset by RBC counts. The health status was determined by the best blood analyte correlate in this case, as previously discussed, the RBC counts, to separate healthy from unhealthy groups by sex. The results show that the immune-score predicted age has the lowest prediction error rate among healthy groups, for both sexes and overestimates age when the individuals are not healthy.
6 Conclusion
We have introduced risk-assessment scores and studied their behaviour against typical synthetic and empirical disease cases. We have also introduced a learning procedure to adapt normal versus abnormal reference values to personalised ranges for more precise and individual medical assessment and demonstrated the capabilities and applications of the scores also for triaging purposes against a popular public database showing how the scores separate healthy from unhealthy people even in the face of significant noise coming from self-reported health status. The numerical score was also shown to be able to define an ‘immune age’, with some significant differences for males and females who are known to have greater fluctuations related to e.g. menopause.
The approaches were shown to be informative for quick and entry-level patient health sorting and triaging purposes. In the case of the immune age, the age of healthy people was very closely predicted mostly by RBC as a component of the CBC, and for unhealthy both by reported or with actual abnormal CBC values, the score was not predictive hence indicating that unhealthy people will find a mismatch between their score age prediction and their actual biological age.
Unlike statistical or traditional machine learning approaches, these scores are transparent so clinicians and health professionals, as well as health consumers, can fully understand how they mechanistically work and allow calculation by hand with each step justified by a mathematical interpretation that can be assessed in a clinical context for rapid screening and decision support.
Appendix
Table 5: List of blood-related label headers in the NHANES database
LBXWBCSI White blood cell count (1000 cells/uL)
LBXLYPCT Lymphocyte percent (%)
LBXMOPCT Monocyte percent (%)
LBXNEPCT Segmented neutrophils percent (%)
LBXEOPCT Eosinophils percent (%)
LBXBAPCT Basophils percent (%)
LBDLYMNO Lymphocyte number
LBDMONO Monocyte number
LBDNENO Segmented neutrophils number
LBDEONO Eosinophils number
LBDBANO Basophils number
LBXRBCSI Red blood cell count (million cells/uL)
LBXHGB Haemoglobin (g/dL)
LBXHCT Haematocrit (%)
LBXMCVSI Mean cell volume (fL)
LBXMCHSI Mean cell haemoglobin (pg)
LBXMC MCHC (g/dL)
LBXRDW Red cell distribution width (%)
LBXPLTSI Platelet count SI (1000 cells/uL)
LBXMPSI Mean platelet volume (fL)
Data
The National Health and Nutrition Examination Survey[28] is a research program conducted by the National Center for Health Statistics (NCHS) of the United States of America. This program offers a public database that contains information relating to the demographic, health, and nutritional status of adults and children in the USA. In particular, it offers individual laboratory results for the thirteen analytes, along with demographic information (such as age, biological sex, among others) for thousands of individuals collected over 20 years.
For the immune age sections, we extracted a subset of 31900 individuals chosen according to the following criteria.
  • 20 years of age or older as stated by RIDAGEYR.
  • The general health condition (all self-reported) variable (HSD010) was stated as Good, Very Good or Excellent.
  • The values for the 13 analytes, along with the general health condition, age and sex (RIAGENDR) were present in the data.